If the points (2,0),(0,1),(4,5)and (0,c)are concyclic, then the value of "c" is ?

Grade:12

3 Answers

Badiuddin askIITians.ismu Expert
13 years ago

Dear tadiparthi

first find the equation of circle passes through point(2,0) ,(0,1) and (4,5)

equation of the circle

x2 +y2 +2gx +2fy +c =0

so put the points

 4+ 4g +c =0 .............1

 1 +2f +c =0 ..............2

39 + 8g + 10f +c =0 ............3

solve these equation

g =-25/12

f =-8/3

c =13/3

so equation of circle

x2 + y2 -25x/6 -16y/3 +13/3 =0

now point(0,c) satisfy this equation

c2 -16c/3 + 13/3 =0

3c2 -16c +13 =0

c=1 ,13/3

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you
the answer and detailed  solution very  quickly.

 We are all IITians and here to help you in your IIT JEE preparation.

All the best.
 
Regards,
Askiitians Experts
Badiuddin

Keshav Jhawar
11 years ago

The third eq is most probably wrong as 5^2 + 4^2 is 41 and not 39 as you have mentioned. If there is any fault in what I said please correct me.

ankit singh
3 years ago
First find the equation of circle passes through point(2,0) ,(0,1) and (4,5)
equation of the circle
x2 +y2 +2gx +2fy +c =0
so put the points
4+ 4g +c =0 .............1
1 +2f +c =0 ..............2
39 + 8g + 10f +c =0 ............3
solve these equation
g =-25/12
f =-8/3
c =13/3
so equation of circle
x2 + y2 -25x/6 -16y/3 +13/3 =0
now point(0,c) satisfy this equation
c2 -16c/3 + 13/3 =0
3c2 -16c +13 =0
c=1 ,13/3